OpenStudy (anonymous):

What is the range of y = 5x -2 if the domain is {-3, -1, 0, 1, 3}? Answer {-13, -3, 2, 3, 13} {12, 3, 0, 3, 13} {-17, -7, -2, 3, 13}

OpenStudy (anonymous):

y = 5x - 2 when x = 0 y = -2 when x = -3 , y = 5*-3 - 2 = -15-2 = -17 when x = 3 y = 5*3 - 2 = 15 - 2 = 13 therefore , answer is the third option